1. Calculate GDP and National Savings from the following data:


  1. Compensation of Employees=7937
  2. Income of self-employed=1006
  3. Rental Income=110
  4. Corporate profits=1260
  5. Indirect business taxes=994
  6. Consumption of fixed capital=1747
  7. Factor income received from rest of the world=709
  8. Payment of factor income to rest of the world=567
  9. Consumption=9765
  10. Government Spending=900
1
Expert's answer
2021-06-29T12:32:56-0400

"GDP = C + G + I + NX"

C  is consumption or every private consumer spending within the economy of a country.

G are all the government expenditures.

I is the sum of investment a country spent towards capital equipment as well as inventory.

NX = Nation's total exports - total imports or net exports.

 

"GDP = 9765 + 1747 + 110 + 1006 + 900""GDP = 13528"